# Optimal alignment of Lorentz orientation and generalization to matrix Lie groups

### v2

@@ IV.1 Ideal conditions with no noise
with $\beta=0.3$ and $\gamma=\frac{1}{\sqrt{1-\beta^{2}}}$, and verified that both methods could reproduce $\Lambda$, given $A$ and $B=A\Lambda^{T}$. This simple sanity check showed that both methods indeed work. The use of the explicit form of the Lorentz boost given by Haber in Eqs (30)-(35) ensured that the resulting matrix indeed had determinant close to $1$. However, the Lie algebra method was slightly more effective than direct least squares optimization via the default solver in scipy.optimize.minimize in ensuring that the components of $A\Lambda_{\text{estimated}}^{T}$ were close to $A\Lambda^{T}$, as demonstrated in the Supplemental Materials. For the Lie algebra method, we tested brute force matrix exponentiation via scipy.linalg.expm compared to Haber’s explicit computation, and found that both computations resulted in $\det\Lambda\approx 1$ to 14 decimal places.
The timing between the two Lie algebra methods was also extremely similar, with each iteration taking 700-800 $\mu$s, with standard deviation of 20-30 $\mu$s. In contrast, the default solver for scipy.optimize.minimize took an average of 24 ms, which was 30$\sim$ 30$\times$ 30 × as slow as the Lie algebra method.
@@ IV.2 Varying the number of vectors given, noise, and averaging over arbitrary Lorentz transformations
Figure 1: The Frobenius ($L^{2}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix. The rows represent varying the noise $\epsilon$ present in the measurement of the vectors in frame $B$. The columns represent varying the number $n$ of vectors upon which the estimation was performed. Evidently, the Lie algebra method and the direct minimization method are equivalent in accuracy across a range of testing conditions.
Note that the joint distribution for $\bm{\zeta}$ and $\bm{\theta}$ is a spherically symmetric product of normal distributions. We generated $n$ random 4-vectors with Minkowski norm $-1$ by sampling $x,y,z\sim\mathcal{N}(0,0.3)$ and setting $t=\sqrt{1+x^{2}+y^{2}+z^{2}}$, and used these vectors as our $v_{A,i}$. We introduced noise in the vector components as measured in frame $B$ by adding random noise $\Delta x,\Delta y,\Delta z\sim\mathcal{N}(0,\epsilon)$ to $x,y,z$, normalizing the time component for $x+\Delta x,y+\Delta y,z+\Delta z$ again, and then boosting by $\Lambda$ to these vectors.
We observe that the overall distribution of both the maximum element-wise error and the overall $L^{2}$ error is quite similar between the two proposed methods (Figs 1, 2).
Figure 2: The max ($L^{\infty}$) norm of the estimated Lorentz matrix from frame $A$ to frame $B$ as compared to the actual Lorentz matrix. The rows and columns represent the same variables as in Fig 1.
In Figs 1and 2, we compare the Frobenius and max norms of the estimated Lorentz transformation matrix against the true Lorentz transformation. We generated $1,000$ random Lorentz transformations using
@@ V Discussion
Instead, we can either perform constrained optimization using gradient descent or related methods (the minimization method) or project the least-squares linear operator via the pseudoinverse onto the Lorentz group via a second least squares minimization in the Lie algebra. We have shown that at least in Python, the second approach is more performant, and better reproduces the answer in the case that an exact solution exists. We did not vary the minimization method used by SciPy’s default solver, though ideally an optimal solution would not be dependent on such as choice. The explicit formula relating the boost and rotation vectors to the corresponding Lorentz transformation was necessary in the minimization method, whereas it was optional in the Lie algebra method with little performance difference between it and brute force matrix exponentiation.
The Lie algebra method presented here technically also solves the orthogonal Procrustes problem by restricting our attention to the bottom right $3\times 3$ matrix. However, it extends to arbitrary matrix Lie groups, unlike the Kabsch algorithm which only works for $O(N)$, and the Horn algorithm which relies on a quaternion representation with bounded components. Therefore we feel that this work is valuable in the overall scheme of manifold alignment. An application of Lorentz alignment is finding the connection between free-falling reference frames in smooth lattice general relativity [11].
Future directions on this topic may provide more extensive benchmarking across other matrix Lie groups, and implement the methods described here in a lower level language.
@@
All code needed to reproduce the data and figures in this work are included in the Supplemental Materials as a Jupyter notebook.
@@ Acknowledgments
The author would like to thank the Penn State College of Medicine’s Medical Scientist Training Program for their support of his career.

### v3

@@ IV.1 Ideal conditions with no noise
with $\beta=0.3$ and $\gamma=\frac{1}{\sqrt{1-\beta^{2}}}$, and verified that both methods could reproduce $\Lambda$, given $A$ and $B=A\Lambda^{T}$. This simple sanity check showed that both methods indeed work. The use of the explicit form of the Lorentz boost given by Haber in Eqs (24)-(30) ensured that the resulting matrix indeed had determinant close to $1$. However, the Lie algebra method was slightly more effective than direct least squares optimization via the default solver in scipy.optimize.minimize in ensuring that the components of $A\Lambda_{\text{est}}^{T}$ were close to $A\Lambda^{T}$, as demonstrated in the attached Code and Data. For the Lie algebra method, we tested brute force matrix exponentiation via scipy.linalg.expm compared to Haber’s explicit computation, and found that both computations resulted in $\det\Lambda\approx 1$ to 14 decimal places.
[…]
@@ IV.2 Varying the number of vectors given, noise, and averaging over arbitrary Lorentz transformations
In Figs 1 and 2, we compare the absolute Frobenius ($L^{2}$) and max ($L^{\infty}$) norms of the estimated Lorentz transformation matrix ($\Lambda_{\text{est}}$) against the true Lorentz transformation ($\Lambda_{\text{GT}}$). We generated $1,000$ random Lorentz transformations using
@@
Figure 1: The absolute Frobenius ($L^{2}$) norm of the estimated Lorentz matrix from frame AA to frame $B$ as compared to the actual Lorentz matrix ($\|\Lambda_{\text{est}}-\Lambda_{\text{GT}}\|_{2}$). The rows represent varying the noise $\epsilon$ present in the measurement of the vectors in frame $B$. The columns represent varying the number $n$ of vectors upon which the estimation was performed. Evidently, the Lie algebra method and the direct minimization method are equivalent in accuracy across a range of testing conditions, since their absolute error distributions largely overlap.
Figure 2: The absolute max ($L^{\infty}$) norm of the estimated Lorentz matrix from frame AA to frame $B$ as compared to the actual Lorentz matrix ($\|\Lambda_{\text{est}}-\Lambda_{\text{GT}}\|_{t}ext{max}$). The rows and columns represent the same variables as in Fig 1.
Note that the joint distribution for $\bm{\zeta}$ and $\bm{\theta}$ is a spherically symmetric product of normal distributions. We generated $n$ random 4-vectors with Minkowski norm 1-1 by sampling $x,y,z\sim\mathcal{N}(0,0.3)$ and setting $t=\sqrt{1+x^{2}+y^{2}+z^{2}}$, and used these vectors as our $v_{A,i}$. We introduced noise in the vector components as measured in frame $B$ by adding random noise $\Delta x,\Delta y,\Delta z\sim\mathcal{N}(0,\epsilon)$ to $x,y,z$, normalizing the time component for $x+\Delta x,y+\Delta y,z+\Delta z$ again, and then boosting by $\Lambda$ to these vectors.
We observe that the overall distribution of both the absolute maximum element-wise error and the overall $L^{2}$ error is quite similar between the two proposed methods (Figs 1, 2).
@@ V Discussion
Instead, we can either perform constrained optimization using gradient descent or related methods (the direct minimization method) or project the least-squares linear operator via the pseudoinverse onto the Lorentz group via a second least squares minimization in the Lie algebra (the Lie algebra method). We have shown that at least in Python, the second approach is more performant, and better reproduces the answer in the case that an exact solution exists. We did not vary the direct minimization method used by SciPy’s default solver, though ideally an optimal solution would not be dependent on such as choice. The explicit formula relating the boost and rotation vectors to the corresponding Lorentz transformation was necessary in the direct minimization method, whereas it was optional in the Lie algebra method with little performance difference between it and brute force matrix exponentiation.
The Lie algebra method presented here technically also solves the orthogonal Procrustes problem by restricting our attention to the bottom right $3\times 3$ matrices in the formulae above. However, it extends to arbitrary matrix Lie groups, unlike the Kabsch algorithm which only works for $O(N)$, and the Horn algorithm which relies on a quaternion representation with bounded components. Therefore we feel that this work is valuable in the overall scheme of manifold alignment. An application of Lorentz alignment is finding the connection between free-falling reference frames in smooth lattice general relativity [13].
The direct minimization method requires repeated evaluation of the matrix exponential in Eqs (24)-(30), on the order of tens to hundreds of times, in order to perform gradient descent or some other method of iterative optimization. In contrast, the Lie algebra method requires a single evaluation of the matrix logarithm which may be computed iteratively [14], and a single evaluation of the matrix exponential. Depending on the convergence speed of the matrix logarithm in comparison to the matrix exponential, the Lie algebra method likely outperforms the direct minimization method for other matrix Lie groups, though further numerical experiments are necessary to verify this hypothesis. In both methods, we used an explicit formula to compute the matrix exponential, however this is also not essential since the matrix exponential can be computed through eigendecomposition. Again, further numerical experiments would be required to compare the performance of explicit formulae for the matrix exponential versus exponentiation via eigendecomposition for other matrix Lie groups. Finally, implementation of these algorithms in other languages such as Julia and C would enable more comprehensive benchmarks.
@@
All code needed to reproduce the data and figures in this work are available at Zenodo (doi: 10.5281/zenodo.15724187).
@@ Acknowledgments
The author would like to thank the Penn State College of Medicine’s Medical Scientist Training Program for its support of his career.
@@
Appendix A Quaternions
Quaternions (often represented by the symbol $\mathbb{H}$) can be thought of as a four-dimensional real vector space with basis $\{\mathbf{1},\mathbf{i},\mathbf{j},\mathbf{k}\}$, with quaternion addition defined by componentwise addition. We will represent quaternions as $q=(q_{0},q_{1},q_{2},q_{3})$, where $q_{0},q_{1},q_{2},q_{3}\in\mathbb{R}$. We denote the imaginary unit in $\mathbb{C}$ by $h^{2}=-1$. The operation of quaternion conjugation is defined by
$q^{*}=(q_{0},-q_{1},-q_{2},-q_{3}),$ (54)
$\mathbf{i}\mathbf{j}=\mathbf{k},$ (55)
$\mathbf{j}\mathbf{k}=\mathbf{i},$ (56)
$\mathbf{k}\mathbf{i}=\mathbf{j},$ (57)
$\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=-\mathbf{1},$ (58)
so that
$(p_{0},p_{1},p_{2},p_{3})$ $(q_{0},q_{1},q_{2},q_{3})$
$=($ $p_{0}q_{0}-p_{1}q_{1}-p_{2}q_{2}-p_{3}q_{3},$
$p_{0}q_{1}+p_{1}q_{0}+p_{2}q_{3}-p_{3}q_{2},$
$p_{0}q_{2}-p_{1}q_{3}+p_{2}q_{0}+p_{3}q_{1},$
$p_{0}q_{3}+p_{1}q_{2}-p_{2}q_{1}+p_{3}q_{0}).$ (59)
The quaternion norm is defined as $q||_{\text{quat}}=qq^{*}=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}$, and unit quaternions by $q||_{\text{quat}}=1$. It is unit quaternions which are isomorphic $SU(2)$, which is a double cover of $SO(3)$. To represent rotations, a vector $\mathbf{r}\in\mathbb{R}^{3}$ is encoded as a quaternion $r=(0,r_{1},r_{2},r_{3})$. We define the operator $U$ which undoes this encoding by $U(r)=\mathbf{r}$. Then the rotation of $\mathbf{r}\in\mathbb{R}^{3}$ by a matrix $R\in SO(3)$ can alternatively be performed in quaternion space by:
$R\mathbf{r}=U(qrq^{*}),$ (61)
where $q$ is the quaternion representing $R$, the left-hand side $R\mathbf{R}$ is a matrix multiplication, and the right-hand side $qrq^{*}$ is a product of three quaternions, computed using Eq (59).
The components of $q$ which correspond to a rotation around an axis $\mathbf{u}$ counterclockwise by an angle $\theta$ are:
$q=\left(\cos\frac{\theta}{2},u_{1}\sin\frac{\theta}{2},u_{2}\sin\frac{\theta}{2},u_{3}\sin\frac{\theta}{2}\right).$ (62)
Using Eq (62) along with Rodrigues’ rotation formula [6], Eq (61) can be directly verified.